OpenStudy (anonymous):

In a right triangle, find the length of the side not given. b=1, c=sqrt26 the length of the third side is ? it ask to simplify, type an exact answer

OpenStudy (anonymous):

5

OpenStudy (anonymous):

just like the last one. you know \[a^2+b^2=c^2\] \[1^2+b^2=\sqrt{26}^2\] \[1+b^2=26\] \[b^2=25\] \[b=5\]

OpenStudy (anonymous):

use the phytagorean theorem, c^2=a^2+b^2 a^2=c^2-b^2 a^2=(square root of 26)^2-1^2 a^2=26-1 a=square root of 25 a=5

OpenStudy (anonymous):

damn sat, you went all out with the TeXing

OpenStudy (anonymous):

$$ \tex $$

OpenStudy (anonymous):

main reason i am here it so learn this. i mean main selfish reason

OpenStudy (anonymous):

good

OpenStudy (anonymous):

how did you do multiline?